### v1

# What Matters in Hierarchical Search for Combinatorial Reasoning Problems?

Abstract
Efficiently tackling combinatorial reasoning problems, particularly the notorious NP-hard tasks, remains a significant challenge for AI research. Recent efforts have sought to enhance planning by incorporating hierarchical high-level search strategies, known as subgoal methods. While promising, their performance against traditional low-level planners is inconsistent, raising questions about their application contexts. In this study, we conduct an in-depth exploration of subgoal-planning methods for combinatorial reasoning. We identify the attributes pivotal for leveraging the advantages of high-level search: hard-to-learn value functions, complex action spaces, presence of dead ends in the environment, or using data collected from diverse experts. We propose a consistent evaluation methodology to achieve meaningful comparisons between methods and reevaluate the state-of-the-art algorithms.
1 Introduction
Many real-world tasks involve underlying abstract structures that capture their complexity. For example, robotic navigation can be efficiently modeled as a graph search problem, enabling optimal path planning and execution within complex environments [37, 9]. Similarly, sequence alignment in bioinformatics can be effectively modeled as a graph of alignment positions and pathways, facilitating the identification of similarity regions [44, 54]. Hence, addressing these problems in discrete environments, which demand intricate reasoning and include numerous NP-hard problems, is crucial for advancing AI [4].
Tackling these challenges requires efficient planning strategies [5]. A recent, promising approach to planning uses subgoals to decompose complex tasks into sequences of simpler steps, drawing inspiration from how humans plan their actions [30, 21, 33]. Hierarchical planning was successfully applied to tasks with high combinatorial complexity by Subgoal Search (kSubS) [14], followed by other algorithms, including Adaptive Subgoal Search (AdaSubS) [63], Hierarchical Imitation Planning with Search (HIPS) [35], and HIPS-$\varepsilon$ [36]. These methods take high-level steps, which are commonly believed to scale better to longer horizons [7, 38], bypass errors in value approximations [14] and reduce complexity by decomposing the problem into subtasks [56, 63].
While subgoal methods have demonstrated strong results in the literature, our findings reveal that classical, low-level search algorithms can perform at least equally well in certain domains. Thus, it remains an open question in which settings subgoal-based methods should be applied to leverage their advantages.
In this work, we present a comprehensive study exploring how the choice of environments and training data impacts subgoal-based methods for combinatorial reasoning problems. To the best of our knowledge, this is the first systematic study to investigate this relationship. We identify key attributes that enable leveraging the advantage of high-level search: hard-to-learn value functions, complex action spaces, the presence of dead ends in the environment, and the use of data collected from diverse experts. Additionally, our experiments reveal several evaluation pitfalls that are easy to overlook but can lead to misleading conclusions. To address these issues, we propose a consistent evaluation methodology that allows meaningful comparisons between various methods.
In summary, our contributions are as follows:
  • We identify the key properties of environments and datasets that impact the performance of subgoal-based methods.
  • We analyze the suitability of Hierarchical Reinforcement Learning (HRL) algorithms for combinatorial reasoning and list the necessary attributes.
  • We highlight critical points for reporting reliable results in this domain, such as reporting success rates for the entire search budget.
2 Related Work
Classical reinforcement learning (RL) algorithms, such as Proximal Policy Optimization (PPO) [49] and Deep Q-Network (DQN) [42], are designed to solve problems modeled as Markov Decision Processes (MDPs) [56]. Recently, there has been a surge in research leveraging large datasets of demonstrations [58, 11, 25, 17] to learn and solve complex real-world decision-making tasks using Imitation Learning (IL) techniques. These methods, including Behavioral Cloning (BC) [55, 48, 58], Inverse Reinforcement Learning [3], and DAgger [47], aim to mimic expert actions provided in datasets.
Subgoal-based methods decompose complex tasks into manageable subgoals, streamlining the solution of complex MDPs. HIRO [43] leverages offline data by relabeling past transitions. HAC [39] learns a hierarchy of high-level policies using hindsight [1]. Graph-based search methods, such as SoRB [18] or DHRL [38], construct a high-level graph of the state space, enabling fast searches for arbitrary goals. Hierarchical Diffuser [7] generates action plans at various timescales using a diffusion model [59]. HIPS [35] employs Policy-guided Heuristic Search [45] to navigate between subgoals generated by VQ-VAE models [57]. HIPS-$\epsilon$ [36] combines high-level and low-level planning to ensure completeness of the search.
Recently, kSubS [14] was introduced as a hierarchical search algorithm for combinatorial domains. To solve the given task, it iteratively constructs a search tree consisting of achievable subgoals. In each iteration, the $k$-subgoal generator predicts states that are $k$ steps closer to the solution, after which the low-level policy computes paths to these subgoals. Nodes are expanded based on a best-first strategy guided by the value function predictions. All components of kSubS are trained on offline datasets using imitation learning.
AdaSubS [63] builds on kSubS by adding multiple subgoal generators trained to output subgoals at various distances. During the planning phase, AdaSubS prioritizes long-distance subgoal generators, retracting to shorter if the search becomes locally stuck. This way, AdaSubS dynamically adjusts its planning strategy based on the difficulty of the current search region.
Classical search algorithms are widely used to solve hard decision-making problems [27, 23, 41, 31, 50, 19, 8]. Best-First Search (BestFS) prioritizes states with the highest predicted values, often using policy to select actions based on top-ranked choices or cumulative confidence thresholds [20]. A* improves upon BestFS by combining value estimates with distances from the root state, enabling more efficient exploration. MCTS [12] builds an asymmetric tree to estimate action values through random simulations, precisely choosing the most promising action. Monte Carlo Tree Search (MCTS) [12] builds an asymmetric tree to estimate action values through random simulations, selecting the most promising actions precisely. Combining MCTS with deep neural networks has led to breakthroughs like AlphaGo and AlphaZero [50, 52].
3 Preliminaries
3.1 Combinatorial environments and algorithmic desiderata
The abstract structure of real-world tasks captures the essence of reasoning. Therefore, our study targets solving discrete combinatorial reasoning environments, a category that includes many NP-hard problems, such as the Traveling Salesman Problem [2], the Knapsack Problem [32], the Rubik’s Cube [53], or the Sudoku [62]. Specifically, in our experiments, we use the Rubik’s Cube, Sokoban, N-Puzzle, and Inequality Theorem Proving (INT) [60]. These are classical benchmarks, explored in various planning works [41, 14], and are known to be NP-hard [15, 13, 46]. Descriptions of these environments are in Appendix A.
The combinatorial reasoning environments are typically characterized by an enormous space of possible configurations, making learning without priors or handcrafted dense rewards prohibitively challenging [^1] For instance, we trained a PPO [49] agent to solve the Rubik’s Cube, but, unsurprisingly, it failed to make any progress due to never reaching the goal in the haystack of $4.3\times 10^{19}$ states, hence never observing a positive reward. This would be the case for any algorithm that learns goal-reaching. Similarly, the space complexity usually limits the number of configurations that can be observed in training. Hence, the solver cannot expect that it will eventually visit a closely located state again, which is essential for methods like [18, 29, 38]. Therefore it follows that the following conditions are necessary for an algorithm to be efficient in solving combinatorial problems:
  1. 1.
    Learning from offline demonstrations Finding a solution through exploration is prohibitively hard in combinatorial environments. Thus, the algorithm has to be able to learn from demonstrations.
  2. 2.
    Generalization The algorithm will typically visit a given state only once, so it needs to generalize well across the state space.
  3. 3.
    Planning The algorithm needs a planning module as it cannot scale in solving NP-hard problems [5] within a fixed computation budget.
Standard hierarchical methods have not been specifically designed for combinatorial problems and, thus, fail to meet the necessary conditions for effective application. For instance, [7, 61] require continuous state or action space, [24] learns from online interactions, [18, 29, 38] assume a good coverage of the whole state space, and [43, 39] do not use planning to determine actions.
3.2 Subgoal methods
Subgoal methods perform planning with high-level steps composed of multiple actions. This high-level guidance may take various forms. We consider subgoal methods characterized by the following components:
  • Subgoal generator that, given a state, outputs subgoals to achieve. That may be a future state, a class of desired outcomes, or any other form of high-level guidance.
  • Low-level policy that determines low-level paths between subgoals. It may be a trained goal-reaching policy [14, 63], a local search [14, 35], a stored path from previous episodes [18, 38].
  • Planner that determines the order of generating subgoals. Usually, standard planners, e.g., BestFS [14], PHS [35], or their modified forms [63], are repurposed for this goal.
  • Value function that estimates the distance to the solution.
In this work, we will study kSubS and AdaSubS as subgoal methods suitable for combinatorial problems. We also considered HIPS, HIPS-$\varepsilon$, but found them impractical (see Appendix I for details). For reporting, we use the notation kSubS-$n$ to indicate that the subgoal generator was trained for distance $n$. Similarly, AdaSubS-$a$+$b$+$c$ indicates using a set of generators trained for distances $a$, $b$, and $c$.
We compare the performance of subgoal search methods with three widely used low-level algorithms: BestFS, A*, and MCTS. Detailed description, analysis, and pseudocode for each of those low-level algorithms are provided in Appendix F.
4 Analysis
In this work, we aim to establish a solid empirical foundation for developing subgoal-based search algorithms tailored to combinatorial reasoning problems. To this end, we pose the following fundamental questions:
  1. Q1.
    Are hierarchical search algorithms indeed proficient in solving combinatorial problems?
  2. Q2.
    In which environments do hierarchical search algorithms offer an advantage over traditional low-level search strategies? Can we identify specific properties that make an environment amenable to high-level search?
  3. Q3.
    How should we select the baselines for a fair comparison? What common mistakes should be avoided when reporting experimental results?
To address these questions, we extensively analyze the algorithms and combinatorial reasoning environments. Each experiment is described starting with its motivation, followed by the presentation of our experimental results, and concluding with concise takeaways. [^2] The code used to run the experiments can be accessed via the following link: github.com/subgoal-verification/what-matters-in-hierarchical-search.
Experimental setup
For each experiment, we generate a dataset of approximately $5\times 10^{5}$ solving trajectories. We train the components for each method through imitation learning. Whenever possible, components are shared between algorithms; for instance, all tested algorithms use the same value function. For clarity, we present only the best-performing variants of each method. Tuning the parameters is discussed in Appendix F.
Our primary metric, success rate, is the percentage of solved instances calculated across a set of $500$ instances in a given environment. The success rate depends on the search budget, which is the total number of states visited during the search, including both the subgoals and the low-level steps.
The costs associated with this metric arise from two main sources: generating subgoals and filling low-level paths, each requiring network calls. Therefore, we argue that reporting all search nodes is essential. For instance, a generator might produce a subgoal that is 10 steps ahead of the current state, but reaching this subgoal could be highly costly due to the associated reaching expenses. However, it would be hidden by the other metric only for subgoal methods. A further discussion of this decision is presented in Section 4.3.1.
4.1 Rubik’s Cube
4.1.1 Verifying experimental outcomes
Motivation
Learning to solve the Rubik’s Cube is hard, even for humans. [14, 63] report that both kSubS and AdaSubS can solve it, while the low-level BestFS struggles. We want to verify these claims and understand the reasons behind them.
Experimental Results
Following [63], each method was trained on trajectories obtained by shuffling the cube with $20$ random moves and reversing it. As shown in Figure 2, every tested method solves the environment nearly perfectly. AdaSubS is the most effective when using smaller search budgets, while BestFS takes the lead for larger budgets. A* performs similarly to BestFS, but MCTS is not effective. Thus, using subgoals in this setting is not helpful. This contrasts with the results presented in [14, 63]. Firstly, the low-level algorithms successfully solve the task. Secondly, each subgoal method requires about three times smaller computational budget than reported in [63].
For this study, we reimplemented and trained the components from scratch, with some technical modifications, such as model size, architecture, and hyperparameters, and using up-to-date libraries. We also noticed that BestFS is very sensitive to the quality of its components and hyperparameters (see Appendix F for details). That may explain the substantial difference in performance.
Takeaway
Low-level search algorithms are competitive with subgoal methods for solving the Rubik’s Cube
4.1.2 Learning from diverse experts
Motivation
Subgoal methods are considered more effective for long-horizon planning problems [7] and are believed to reduce noise in training data [14]. To verify these properties, we trained the algorithms to solve the Rubik’s Cube using a dataset composed of various sources: human-oriented Beginner and CFOP, computational Kociemba, and the random solutions used in Section 4.1.1. The first two solvers solve the cube layer by layer; hence, their trajectories are long and composed of high-level subroutines. The remaining two usually find much shorter solutions (see Appendix B.2 for examples).
Experimental Results
When training on data from diverse experts, the subgoal methods (particularly AdaSubS) outperform the low-level search by a wide margin (see Figure 2). Notably, despite using strongly suboptimal and diverse data, AdaSubS and kSubS can solve the most challenging instances within an even smaller budget than in the previous experiment. On the other hand, all low-level methods perform much worse than in the previous experiment. Additionally, we observe that given lengthy demonstrations, increasing the subgoal distance for kSubS improves performance, unlike in Section 4.1.1 with short trajectories.
We attribute the observed performance gap to the noise in the value function estimates. Specifically, it fails to consistently assign low values to fully scrambled cubes and nearby states. This issue arises because both Beginner and CFOP yield solutions that are even five times longer than those of the other solvers. Subgoal algorithms are, to some degree, robust to this noise, as claimed by [14]. By taking longer steps, they can leap over the faulty value function. A detailed analysis of this behavior is provided in Appendix B.2.
Takeaway
Subgoal methods successfully leverage diverse, long demonstrations, while low-level search performs better on short, homogeneous trajectories.
Figure 1: Solving the Rubik’s Cube. Components are trained on reversed random shuffles.
Figure 2: Solving the Rubik’s Cube. Components are trained on data from 4 different experts.
4.1.3 Structure in demonstrations
Motivation
While Section 4.1.2 showcases an advantage of subgoal methods, it is still unclear which property of the environment contributes the most to it. We conjecture that the key factor is the diversity of training data, which makes the value estimation problem hard. To verify this, we use training data collected from a single algorithmic solver – Beginner. Unlike in the previous experiment, no other expert is used. Therefore, the training data, although not optimal, becomes unambiguous.
Figure 3: Solving the Rubik’s Cube. Components are trained on the Beginner algorithmic solver.
Figure 4: Success rate of solving the Rubik’s Cube within a budget of $500$ visited states, with artificial Gaussian noise added to each value estimate.
Experimental Results
After removing the variance from demonstrations by switching to Beginner, the performance of AdaSubS and BestFS becomes nearly identical, as illustrated in Figure 4. Similarly, the performance of A* closely matches that of kSubS. MCTS shows significant improvement over previous experiments, though it still lags behind other methods. Longer subgoals continue to outperform shorter ones, but extending them beyond a distance of four offers no additional benefits. Furthermore, a comparison with the results in Section 4.1.2 reveals that the performance of subgoal methods remains consistent whether using single or multiple sources of demonstrations.
Importantly, we found that in order to achieve good performance, BestFS requires careful tuning. It achieves its best results when, at each expansion, we select actions that collectively exceed $70\%$ confidence. Additionally, we found that its performance is highly sensitive to this hyperparameter (see Appendix F). Adopting the approach recommended by [14, 63] of selecting a fixed number of actions leads to highly suboptimal outcomes, as demonstrated by the poorly performing agent in Figure 4, BestFS-top-2.
Takeaway
Given homogeneous data, low-level search algorithms can leverage the long, structured demonstrations equally well as subgoal methods.
4.1.4 Robustness to noisy value function
Motivation
The advantage of subgoal methods observed in Section 4.1.2 does not primarily come from long or structured demonstrations. However, high-level algorithms show surprisingly consistent performance regardless of the training data, as demonstrated in Sections 4.1.1-4.1.3. That observation hints that the subgoal methods are resilient to noise in the training data. To verify that proposition, we take the components trained as in Section 4.1.1 and add artificial Gaussian noise to each output of the value function.
Experimental Results
As shown in Figure 4, while the standard BestFS can solve nearly all instances, its performance heavily depends on the quality of the learned value function. In a similar way, the noisy values strongly impact the performance of A* and MCTS. However, kSubS is much more resilient when using subgoals longer than a single step. Even with completely random values ($\sigma=100$), it successfully solves about $40\%$ instances. Moreover, AdaSubS is barely affected by the noise at all. That confirms the claim of [14] that subgoals successfully leap over value noise.
As noted in [63], the search performed by the subgoal methods is strongly guided by the subgoal generators, which considerably reduces the reliance on the value function. A more detailed analysis of this experiment can be found in Appendix B.3.
Takeaway
Subgoal methods successfully leap over the noise. Thus, they should be used when estimating the value is hard, for instance, when learning from diverse and suboptimal experts.
4.2 INT
4.2.1 Scaling to complex mathematical reasoning
Motivation
We aim to explore whether subgoal methods can more effectively scale to complex real-world applications than low-level methods, as suggested in [14, 63]. To this end, we consider the INT environment [60], which involves solving inequalities. It is the most challenging among the environments studied in [14, 63, 35, 36], both in terms of observations and actions complexity.
Experimental Results
Our experiments show a clear advantage of subgoal methods compared to low-level search on INT, as shown in Figure 6. We note that we use the proofs length of $25$, which is a more challenging setup than typically used.
Takeaway
Subgoal methods significantly outperform low-level search strategies in the context of mathematical reasoning.
4.2.2 Complex action space
Motivation
We conjecture that the result of the previous experiment is due to the high complexity of the action space compared to the state space. This suggests that state-based subgoals may be more effective in such settings than action-based policy guidance. To further investigate this in a controlled setting, we inflate the action space of the Rubik’s Cube by a factor of $100$ by creating $100$ copies of each original action.
Experimental Results
As shown in Figure 6, the low-level search methods stuck below $30\%$ success rate, while the subgoal methods perform comparably to the standard cube case from Section 4.1.1. We could expect that since, given so many similar actions, the low-level search nearly trivializes. Although the subgoal methods also use a low-level policy for connecting the subgoals, it always follows a single path towards the subgoal, so it is barely affected. A more detailed analysis can be found in Appendix B.4.
Takeaway
When facing a problem with a complex action space, subgoal methods should outperform low-level search.
Figure 5: Solving INT. Components are trained on randomly generated proofs of length $25$.
Figure 6: Solving the Rubik’s cube with inflated action space.
4.3 Sokoban
4.3.1 Importance of budget measures
Motivation
Given that Sokoban is PSPACE-complete [13], it presents a challenging testbed for AI planning algorithms. [63] reports that AdaSubS outperforms BestFS in Sokoban. They also report that kSubS surpasses BestFS only, given a relatively high search budget. Some works use the number of high-level nodes visited during the search as the search budget [35]. This contrasts with the total number of visited nodes (the main metric in this paper). We aim to show how this difference affects the interpretation of the results and propose recommendations for reporting the search budget. For these experiments, each method was trained on a dataset containing trajectories collected by an external RL agent.
Experimental Results
Figure 8 shows that reporting only the high-level nodes excessively enhances subgoal methods’ scores. AdaSubS and kSubS seem absolutely superior to BestFS despite the last two algorithms visiting a similar number of states during the search. Hence, that way of reporting the results would lead to erroneous conclusions about the strong primacy of subgoal methods, and we recommend using the total number of visited states as the search budget.
Takeaway
When reporting the search budget, it is crucial to count all visited states, not only the high-level nodes.
Search Algorithm Dead-Ends
MCTS 22.0%
BestFS-$85\%$ 18.5%
A* 13.7%
kSubS-$4$ 12.7%
kSubS-$8$ 10.0%
AdaSubS-$8+4$ 8.86%
Figure 7: Comparison of the fraction of dead-ends encountered during the problem-solving process between subgoal-based and low-level searches.
Figure 8: Comparison of two meanings of search budget: total states visited versus only high-level nodes visited. For BestFS, both definitions are equivalent.
4.3.2 Dead-ends
Motivation
We call a state a dead-end if it is impossible to reach the goal from it. They pose an additional challenge as an algorithm must quickly recognize dead-ends to abandon the search in hopeless state-space regions. We investigate how the subgoal and low-level methods handle dead-ends.
Experimental Results
Table 8 indicates that subgoal methods, especially those involving longer subgoals, notably reduce the likelihood of stacking in dead-ends. We conjecture that this effectiveness stems mainly from relying less on value function guidance than low-level methods, as detailed in Section 4.1.4 and supported by [20]. For further analysis, refer to Appendix B.1.
Takeaway
Subgoal methods are better in avoiding dead-ends than low-level search.
4.4 N-Puzzle
Motivation
The N-puzzle, also known as the Sliding Tile Puzzle, is a classic benchmark in the search literature [41, 35, 34] and is known to be NP-hard [46]. It was used to evaluate subgoal methods in [35, 36]. Despite its large state space, the N-puzzle is highly structured, allowing for solving very simple suboptimal strategies. We aim to determine whether subgoal techniques offer an advantage over low-level search methods in this environment.
Experimental Results
To generate the training dataset for this experiment, we employed a sub-optimal algorithm that orders the tiles in ascending order, starting with "$1$" and continuing through "$2$", "$3$", and so on (details in Appendix A).
Figure 9: Solving the N-Puzzle. The success rates of BestFS, A*, and BC policy throughout all budgets match nearly perfectly.
A simple policy trained with behavioral cloning (BC) perfectly solves the environment, even without using search, see Figure 9. This indicates that learning from structured demonstrations is relatively straightforward. Notably, AdaSubS surpassed BC, indicating that this subgoal approach improved the cloned sub-optimal behavior. Furthermore, kSubS, BestFS, and A* perform similarly to the BC policy, showcasing no computational waste.
Takeaway
Due to its structure, the N-Puzzle environment is not a good benchmark for testing the performance of advanced search algorithms.
5 Open questions and future directions
While we identified several features that facilitate the performance of subgoal methods, that list is not exhaustive. Thus, it is essential to study this topic further, expand the analysis to more subgoal-based and low-level algorithms, and include even more types of environments. Additionally, our work provides mostly experimental validation of the claims. Finding theoretical foundations for the observed properties would also be a valuable direction.
6 Conclusions
We conducted a thorough comparison of subgoal-based and low-level search algorithms for combinatorial reasoning tasks. Our experiments indicate that hierarchical approaches excel in environments where value estimation is challenging and learned estimates face significant uncertainty, particularly when learning from various suboptimal experts. Furthermore, subgoal methods demonstrate better scalability in complex action spaces and are more effective at avoiding dead ends than low-level methods. Thus, in environments characterized by those properties, it is advisable to consider subgoal-based methods.
According to our experiments, the best-performing low-level search was usually BestFS with a confidence threshold. Although it is rather sensitive to the threshold value, which has to be optimized for each domain separately, we advocate using it as a standard baseline for further research in hierarchical search.
Finally, we identified a few easy-to-overlook mistakes in reporting the results that may lead to misleading conclusions. Most importantly, the reported search budget of the hierarchical methods must include all the visited states and not only the high-level nodes as used in some prior works.
Appendix A Environments
Sokoban Sokoban is a classic puzzle game where the objective is to push boxes onto target locations within a confined space. It is a popular testing ground for classical planning methods and deep-learning approaches due to its combinatorial complexity and difficulty in finding solutions. Recognized as a PSPACE-hard problem, Sokoban is used to evaluate different computational strategies. Our experiments use 12 × 12 Sokoban boards with four boxes to assess the performance of our proposed models. An illustrative example of a simple Sokoban search tree with a solving path is shown in Figure 10.
Figure 10: Hierarchical Search applied to solving Sokoban. This tree, depicted in figures, employs bolded green arrows to highlight selected subgoals within a hierarchical search framework earmarked for subsequent exploration. The illustration demonstrates that these intermediate goals exhibit variability in terms of both their spatial distance and the methodology by which a planning algorithm may leverage them.
Rubik’s Cube The Rubik’s Cube, a renowned 3D puzzle, has over $4.3\times 10^{19}$ possible configurations, highlighting the huge search space and the computational challenge it poses. Recent advancements in solving the Rubik’s Cube with neural networks underscore the potential of deep learning methods in navigating complex, high-dimensional puzzles. For the exact representation of the Rubik’s Cube state, see Figure 11.
wbrwyggwwoboybygbryrorroboygrbggbggbwybrooogrywrowywwy $s_{0}$ Initial State wbrwyggggobwybwgbgooyrroyrbrrbwgbygbwybrooogroyrowywwy $s_{1}$ One Action (= single rotation) wbywyoggbobwybwgbgoorrryyryywrggrbbbgybgoorgroyoowrwww $s_{2}$ gyowyoggbwbwwbwwbgoorrryyryywwggbbbyboryogggroyoowrbrr $s_{3}$$\dots$ yyyyyyyyybbbbbbrrrrrrrrrgggggggggooooooooobbbwwwwwwwww $s_{n-1}$ yyyyyyyyybbbbbbbbbrrrrrrrrrgggggggggooooooooowwwwwwwww $s_{n}$ Solving State
Figure 11: Example trajectory of Rubik starting from initial state $s_{0}$ leading to the final solution $s_{n}$.
N-Puzzle The N-Puzzle, a classic sliding puzzle game, comes in various sizes, including the 3x3 (8-puzzle), 4x4 (15-puzzle), and 5x5 (24-puzzle). The goal is to rearrange a frame of numbered square tiles into a specific pattern, a task that tests the algorithm’s ability to plan and execute a sequence of moves efficiently. Figure 12 shows a visualization of a trajectory in 24-puzzle.
Figure 12: Example trajectory of n-puzzle starting from initial state $s_{0}$ leading to the final solution $s_{n}$. Red arrows indicate low-level actions.
INT INT (INequality Theorem proving) is an automated theorem-proving benchmark for high school algebraic inequality proofs. [60] provides a generator of mathematical inequalities and a proof verification tool. Each action in INT maps to a proof step, which specifies a chosen axiom and its input entities - which makes action space very high-dimensional, enabling up to a million valid actions at a step. This large action space makes INT a desirable but challenging environment for expanding HRL paradigms to vast action spaces.
We used 25-step proofs for this paper, representing an uplift from 15 considered in [14, 63] (the latter used longer proofs, but only for evaluating 15-trained models). Each step is an application of an axiom to an axiom-specific number of entities (entities are bracketed or bracketable parts of the theorem’s goal).
[…]
Figure 13: A comprehensive representation of theorems pertaining to goal achievement in mathematical expressions, showcasing the logical structure and underlying premises leading to the formulated goals.
Appendix B Key factors for Hierarchical Search
According to our experiments, the attributes pivotal for leveraging the advantages of high-level search include:
  • presence of dead ends,
  • learning from diverse experts,
  • hard-to-learn value function,
  • complex action space
In Section 4, we show our main experiments that support our findings. In this appendix, we present an extended analysis of each property.
B.1 Dead ends
In this paper, we conceptualize the environment $\mathcal{E}$ as a Markov Decision Process (MDP), where the environment, for any given state $s\in\mathcal{S}$ and action $a\in\mathcal{A}$, transitions to a subsequent state $s^{\prime}\in\mathcal{S}$. Within this framework, we define a dead-end state $s_{d}$ as a specific type of state with two defining characteristics:
  1. 1.
    The state $s_{d}$ does not represent a solution state.
  2. 2.
    From state $s_{d}$, there exists no feasible sequence of actions leading to a trajectory $(s_{d},s_{d+1},s_{d+2},\dots,s_{d+n})$ that culminates in a solution state $s_{d+n}$.
Dead-end states pose significant challenges in reinforcement learning, as they represent points from which no further learning or progress toward a solution can be made. Identifying and addressing these states is crucial for improving the efficiency and effectiveness of learning algorithms in navigating complex state spaces. Figure 14 shows an illustrative example of a dead-end state.
Figure 14: An example dead-end in Sokoban – a box that is pushed to the corner cannot be moved anymore, so the objective is not possible to achieve.
Examples of dead-ends in kSubS vs. BestFS
In this subsection, we present examples of how each method handles dead-end situations during the search process.
For this presentation, we analyzed 128 search trees initiated from identical starting boards for both algorithms. The kSubS algorithm encountered dead-ends in 3 instances. To resolve these, it navigated through 13 high-level nodes and 105 low-level nodes within the corresponding subtrees. In contrast, the BestFS algorithm encountered dead-ends in 18 instances, requiring the traversal of 4431 nodes. Note that BestFS does not distinguish between high-level and low-level nodes in its search.
Examples of dead-end handling are shown in Figure 15 for kSubS and Figure 16 for BestFS. Observe that in the case showed in Figure 15 expanding the parent node resulted in adding two more dead-ends to the search tree. Because they have higher values, they were immediately expanded. However, the subgoal generator understood that the only way to reach solution is to make an invalid transition of releasing the blocked box. Such subgoals cannot be achieved by the conditional policy, hence no more subgoal was created in that branch. On the other hand, low-level search is unable to propose invalid transitions, so it stays in dead-end until the value estimates are higher than for other branches.
Figure 15: We illustrate a scenario where the kSubS algorithm encounters dead-ends, hindering the search process. The figure shows a case where the algorithm generates two subgoals at an expected distance (k=8), but both lead to dead-ends, wasting a portion of the search budget (18 nodes). As a result, the kSubS algorithm backtracks from this subtree and continues searching elsewhere within the tree.
Figure 16: The figure shows BestFS expanding two nodes from a dead-end. This resulted in the exploration of over 300 additional nodes from that state, ultimately failing to find a solution within the given search budget.
B.2 Learning from diverse experts
Usually broad datasets of offline demonstrations contain data from various sources [10, 26, 28]. Because of that, the demonstrations may be diverse in terms of length, structure, quality, skills, strategy, etc. We want to answer the question whether such setting is handled better by high-level or low-level search algorithms.
Experiment setup For this experiment, we choose the Rubik’s cube environment. We collected a dataset of 500 000500000500\,000500 000 trajectories, computed with four different solvers for the Rubik’s cube:
  • Beginner – the simplest human-oriented solving algorithm. It aims to order the cube layer by layer with a few primitive tactics. Because of that the solutions are structured, but also very long (typically between 150 and 200 moves).
  • CFOP – an algorithm designed for speedcubers. It is based on the same principle as Beginner, but employs many advanced tactics that make the solutions faster (typically about 100 moves).
  • Kociemba – a computational solver that finds near-optimal solutions (usually between 20 and 40 moves) in short time. It is heavily optimized based on the algebraic properties of the Rubik’s cube.
  • Random – solutions obtained by scrambling an ordered cube with random moves and reversing the trajectory.
Figure 27 shows example solutions generated with each solver. Clearly, the algorithmic solvers (Beginner and CFOP) generate much longer solutions that the other methods. They are also more structured, as they are based on building patterns. The computational solver Kociemba on the other hand go directly towards the solution because its moves are carefully optimized to ensure maximal advantage. Because of that, this dataset represent a truly diverse set of demonstrations.
Results As shown in Figure 2, the subgoal methods outperform the low-level methods by a wide margin. While BestFS is comparable on small budgets, it struggles with solving most of the instances. Also, it should be noted that the performance of the subgoal methods changes only slightly compared to training on a single Random solver (Figure 2) while the low-level searches are heavily affected.
Learned values To find the sources of that outcome, we checked the values learned by the heuristic function. Because of the diversity introduced by combining the experts, we should expect that the estimates are subject to high uncertainty and possibly high variance.
Figure 17 shows the distribution of the learned heuristic for random fully shuffled cubes. Although most instances can be solved optimally within 20-26 moves, the estimates range from 14 to 90 steps. Furthermore, the distribution is clearly bimodal – one mode correspond to a typical length of Kociemba solution, the other to CFOP.
Figure 17: Value distribution for fully scrambled cubes, learned on data coming from diverse experts. The values are rescaled so that the x-axis represent the estimated number of steps to the solution. The values represent the mean of each interval.
Furthermore, Figure 22 shows the distribution of value estimates throughout the solutions for each solver. We observe that for the algorithmic solvers the initial distance is considerably underestimated. After about $20\%$ moves the value network recognizes the pattern of layers built by the solvers and expect a long solution by assigning values close to $100$. On the other hand, the values learned for the states visited by the computational solvers start as overestimated, but steadily decrease towards 0.
Figure 18: Beginner solver
Figure 19: CFOP solver
Figure 20: Kociemba solver
Figure 21: Reversed random 20-move trajectories
Figure 22: The learned value estimates distribution for various solvers. For each plot 100 episodes were solved using the respective solver. The boxes represent the distribution of value estimates for the consecutive points of the solution. The x-axis denotes the relative part of the trajectory (i.e., $0.5$ denotes the middle point in each trajectory, regardless of its length). The blue line indicates the true number of steps to the solution.
While it is a reasonable strategy for the value to fit to the provided dataset, it creates a challenge for the search. If a search algorithm aims to imitate Beginner or CFOP, it has to reach the layer pattern, characteristic of those solvers. However, the random states tend to have very low distance estimate, compared to the initial layer patterns. Because of that, for tens of steps the heuristic estimates would be actually increasing, making the reached states less and less probable to expand.
In practice, the low-level searches usually fail to cross this gap. On the other hand, the high-level methods are partially guided by the subgoal generators that ignore the values. The value gap that spans across about 30 steps can be crossed with as few as 5 subgoals of length 6. Because of that both kSubS and AdaSubS can successfully leverage the schematic algorithmic solutions.
To finally confirm that conclusion, we must answer the question whether the performance of low-level searches would increase if they could leverage the algorithmic solutions as well. For that purpose, we trained the components for each method using data only from the Beginner solver. This way we remove the challenge of noisy initial values. As shown in Figure 4, the low-level searches indeed perform much better. BestFS even matches the performance of AdaSubS. That confirms our observation that low-level searchas fail to utilize multimodal data because they rely too much on the value function and seek monotonic slopes.
At the same time we observe that since BestFS and AdaSubS show nearly identical performance on Beginner solutions, it is questionable that hierarchical methods handle long-horizon tasks better, which is a common belief [43, 18, 7].
Figure 23: Beginner
Figure 24: CFOP
Figure 25: Kociemba
ybrbyygyybywwboybyoggwrrgoroobwgogrowwobogbgbrgwrwyrrw ybbbybgywbywwboybyogrwrygoybooogrowgwwoyogwgbrggrwrrrr ybrbyygyybywwboybyoggwrrgoroobwgogrowwobogbgbrgwrwyrrw boobyygyyrywbboybyoggwrrgoroowwgrgrrbbwgowbgorgwrwybwy oyyoyybbgbbwbboybyrywwrrgoroggwgrgrroowgowbgorgwrwybwy oyyoyyowgbbgbbbybbwrryrorwgwggggrrrroowgowbgowoyrwybwy yygyywooooowbbbybbbbgyrorwgwrrggrrrrwgggowbgowoyrwybwy oygwywgooybobbobbwybgyroowgwrrggrrrrwgbgorbgwboyywyrwy yygyywooooowbbbybbbbgyrorwgwrrggrrrrwgggowbgowoyrwybwy yygyyooogoowbbbybbbbyyryrwyrgwrgrrrroggwowggowobrwgbww yygyyooogoowbbbggobbyyryybbrgwrgrrwyoggwowrrrbrwwwowgb yyyyyyooboowbbbggobbwyroybbrrrwggyrwgggoowgrrbrrwwwwgo yyyyyyooboowbbbgrrbbwyroggorrrwggybbgggoowyrwwwbgwrowr yyyyyyrbwoowbbwgrbgybgrbooworroggbbbgggoowyrwywrgwrowr yyyyyyrbwoowbbwyrwgybgrbgrborroggoowgggoowbbbogywwwrrr yyyyyyooooowbbbyrrbbbyrrgggyrrgggoowgggoowbbbwwwwwwrrr yyyyyyooooowbbbbbbbbbyrryrryrrgggggggggoowoowrwwrwwrww oyyoyyoyybbbbbbbbbyrryrryrrgggggggggoowoowoowrwwrwwrww yyyyyyyyybbbbbbbbbrrrrrrrrrgggggggggooooooooowwwwwwwww yyoyyoyyobbbbbbbbbrryrryrrygggggggggwoowoowoowwrwwrwwr yyyyyyyyybbbbbbbbbrrrrrrrrrgggggggggooooooooowwwwwwwww
Figure 26: Random
Figure 27: Example solutions computed by each solver. Because the algorithmic solvers typically require over $100$ steps, we use a tiny font to display it.
B.3 Noisy value function
Accurate learning of value function presents a significant challenge in offline reinforcement learning, where the agent’s training is conducted on a fixed dataset of pre-collected trajectories. Section B.2 hints that when value estimates are subject to high uncertainty, subgoal methods should outperform low-level searches.
To confirm that intuition, we run an experiment in a Rubik’s cube environment (Section 4.1.4). This time, the training setup is the same as in Section 4.1.1. However, during inference, we add additional noise to the value estimates. That is, whenever a node is added to the search tree and its value estimate equals $\hat{v}$, we add it with the value of $\hat{v}+\mathcal{N}(0,\sigma)$ instead.
Figure 4 shows that as the amount of noise increases, each low-level method gets less and less efficient. On the extreme, when using fully random values ($\sigma=100$), they struggle to solve any instance.
On the other hand, subgoal methods are much more resilient to noise in the value. Adaptive Subgoal Search is nearly not affected by the presence of noise. For subgoal distances longer than 1, kSubS retains as much as $40\%$ success rate, even with completely random values. The kSubS-1 variant behaves similarly to BestFS, for they are, in fact, equivalent.
Observe that the search performed by low-level methods is guided mainly by the value function. Hence, if the computed estimates are subject to high variance, low-level search struggles to make any progress. On the other hand, the subgoal search is guided both by the value function and the subgoal generator. Both the subgoal generator and the conditional policy that connects subgoals do not depend on the values. Hence, the value function is used only in the high-level nodes, which is only a fraction of the search tree.
An extreme case of that behavior is demonstrated by Adaptive Subgoal Search. Because in our configuration each generator outputs a single subgoal, the value is nearly not used at all for search. Only when the search is stuck, the secondary generators select the highest-ranked node to expand, which in this case is simply a random node of the tree. To summarize, given random value estimates, AdaSubS reduces to the following strategy:
  1. 1.
    Start from the root node,
  2. 2.
    Move from the current node to the subgoal until possible,
  3. 3.
    If the search is stuck, expand a random node in the search tree with a secondary generator and return to (2).
The experiments show that this simple strategy is surprisingly competitive to the greedy best-first approach, even without noise. Interestingly, it could be implemented in low-level search as well. We leave that promising experiment for future work.
B.4 Complex action spaces
The primary difference between low-level methods and subgoal methods is that the former predicts the next action, and the latter – the next state. In many environments, the action space is as simple as a few bits, allowing for iterating over all possible actions, and sampling them. At the same time, states may be considerably larger, up to the extreme of image observations. However, in some environments, the action space is comparable to the state space, or even more complex. A classic example is the AntMaze environment, in which actions are 8-dimensional, while the goal space is only 2-dimensional [22].
Among the combinatorial reasoning environments we consider, INT has the most complex action space. In INT, actions correspond to proof steps and are represented as the chosen axiom, specification of its input entities, and the required premises [60]. Thus, the complexity of the action is at least comparable to the states. Moreover, solving the INT inequalities is based on constant simplification of the given expression, so the state is getting even smaller with each step.
Our experiments, shown in Figure 6, clearly confirm the advantage of using subgoal methods in the INT environment. To further verify the source of that advantage, we conducted another experiment, in a modified Rubik’s cube environment. Recall that the experiment presented in Section 4.1.1 shows that subgoals offer no significant advantage in the original Rubik’s cube. Now, we want to check whether the outcome would be different if the action space were more complex. For that purpose, we extended the action space 100 times. That is, the new action space consists of 1200 possible moves to choose from – 100 copies of each original action.
As shown in Figure 6, the subgoal methods are barely affected by the change, while the low-level searches are unable to exceed $20\%$ success rate. That result confirms our proposition that when facing a complex action space, hierarchical methods offer considerably better performance.
According to our analysis, the primary issue with low-level searches in the augmented Rubik’s cube is the lack of diversity of visited states. When for each state there are hundreds of actions that lead to a similar outcome, they are rated similarly by the policy. Hence, all the top actions essentially lead to the same outcome, which strongly limits the branching factor and trivializes the search trees. On the other hand, subgoal methods are not affected because subgoal generation does not depend on the action space. The conditional policy that connects the generated subgoals does not build a search tree, but always follows the single best action. Because of that, subgoal methods maintain their performance, even though the action space is much more complex.
It is also important to note that even though some state spaces may seem complex, the underlying manifold of possible configurations is in fact low-dimensional. For instance, we use 12x12 Sokoban boards, where each square is encoded as one-hot of 7 possible items, so technically the state space is 1008-dimensional, while there are only 4 actions. However, in practice the subgoal is defined by the positions of agent and boxes, which is at most 10-dimensional, hence rather simple to generate.
Appendix C Network architectures & training details
We used BART [40] and BERT [16] architectures from HuggingFace Transformers for all components. Subgoal generators and INT’s policies (CLLP and baseline policy) use BART. The remaining policies and value functions use BERT. Following the practice in [63], we’ve reduced model size parameters, as detailed in Table 2.
INT As states in INT are complex objects, we prefer to use their string representations and avoid mapping arbitrarily generated strings into complex states. Requisite modifications to the component definition are best illustrated analogously to Section D.1. A generator is redefined as follows:
$\mathcal{G}_{\text{int}}:\underbrace{\mathcal{S}}_{\text{state to expand}}\rightarrow\underbrace{P(\mathcal{T})}_{\text{set of {proposed} subgoals (in string format)}}$
and conditional level policy:
$\mathcal{P}_{\text{int}}:\underbrace{\mathcal{S}}_{\text{current state}}\times\underbrace{\mathcal{T}}_{\text{subgoal {representation}}}\rightarrow\underbrace{\mathcal{A}}_{\text{action}}$
Sokoban Unlike prior work [63, 14], which used convolutional networks for all components, we work on tokenized representations of Sokoban boards and use BERT/BART architectures instead. This modification did not adversely impact our ability to replicate AdaSubS and kSubS results.
Training pipeline We trained our models from scratch using the HuggingFace Transformer pipeline. Detailed training parameters, which varied across environments, can be found in Table 1.
Infrastructure For training, we used a single NVIDIA A100 40GB GPU node, and each component’s training took up to 48 hours. Because we used pre-trained trajectories, we did not need to use more than one core during training. We ran an evaluation using 24-core CPU jobs on Xeon Platinum 8268 nodes with 192GB of memory.
Environment Hyperparameter Generator CLLP Value Policy
INT learning rate 0.0001 0.0001 0.0003 0.0001
learning rate scheduling linear linear linear linear
warmup steps 4000 4000 2000 4000
batch size 32 32 128 32
weight decay 1e-05 1e-05 1e-05 1e-05
dropout 0.1 0.1 0 0.1
Rubik’s Cube learning rate 0.0001 0.0005 3e-7 0.0001
learning rate scheduling linear linear linear linear
warmup steps 5000 50000 50000 1000
batch size 512 5000 5000 2048
weight decay 0.0001 0.001 0.00001 0.0001
dropout 0.1 0 0 0
Sokoban learning rate 0.00001 0.0001 0.0001 0.0001
learning rate scheduling linear linear linear linear
warmup steps 2500 1000 1000 1000
batch size 512 2048 2048 2048
weight decay 0.0001 0.0001 0.0001 0.000001
dropout 0 0.1 0 0
N-Puzzle learning rate 0.0001 0.0001 0.0001 0.0001
learning rate scheduling linear linear linear linear
warmup steps 5000 2000 2000 2000
batch size 4096 4096 512 4096
weight decay 0.00001 0.00001 0.00001 0.0001
dropout 0.1 0 0 0
Table 1: Training-related hyperparameter values
Environment Hyperparameter Generator CLLP Value Policy
INT d model 512 512 - 512
decoder layers 6 6 - 6
intermediate size - - 256 -
encoder attention heads 8 8 - 8
hidden size - - 128 -
num hidden layers - - 2 -
decoder ffn dim 2048 2048 - 2048
encoder ffn dim 2048 2048 - 2048
encoder layers 6 6 - 6
decoder attention heads 8 8 - 8
Sokoban d model 256 - - -
decoder layers 3 - - -
intermediate size - 512 128 512
encoder attention heads 4 - - -
hidden size - 512 128 512
num hidden layers - 6 1 6
encoder ffn dim 2048 - - -
decoder ffn dim 1024 - - -
encoder layers 3 - - -
decoder attention heads 4 - - -
N-Puzzle d model 64 - - -
decoder layers 3 - - -
intermediate size - 128 128 256
encoder attention heads 4 - - -
hidden size - 128 128 256
num hidden layers - 2 1 3
encoder ffn dim 64 - - -
decoder ffn dim 64 - - -
encoder layers 3 - - -
decoder attention heads 4 - - -
Rubik’s Cube d model 256 - - -
decoder layers 3 - - -
intermediate size - 512 128 512
encoder attention heads 4 - - -
hidden size - 512 128 512
num hidden layers - 2 1 6
encoder ffn dim 2048 - - -
decoder ffn dim 1024 - - -
encoder layers 3 - - -
decoder attention heads 4 - - -
Table 2: Model-related hyperparameter values
Appendix D Offline pretraining
Models are pretrained using an offline imitation learning approach. Specifically, given a set of solution trajectories $\{(s_{0},s_{1},\dots,s_{n_{i}})\}_{i=1}^{N}$ produced by an expert $\mathcal{M}$, or multiple experts $\{\mathcal{M}_{j}\}_{j=1}^{M}$ in cases where offline trajectories are collected from multiple experts, the objective is to learn from these trajectories. It is important to note that these trajectories are not required to be optimal; they may include loops or numerous redundant actions. Description of all components can be found in section D.1 and supervised training objectives in section D.2.
D.1 Components
During the pretraining phase, models undergo an offline imitation learning process. Specifically, they are trained on a set of solution trajectories $\{(s_{0},s_{1},\dots,s_{n_{i}})\}_{i=1}^{N}$, which are collected to facilitate the learning of decision-making strategies.
Generator The generator component is responsible for generating subgoal propositions upon receiving a state. These propositions are designed to facilitate progress toward the solution by suggesting intermediate steps that direct the search process more efficiently.
$\mathcal{G}:\underbrace{\mathcal{S}}_{\text{state to expand}}\rightarrow\underbrace{P(\mathcal{S})}_{\text{set of subgoal propositions}}$
Conditional Low-Level Policy The Conditional Low-Level Policy (CLLP) plays a crucial role in node expansion by evaluating each subgoal proposition. For a given current state and a subgoal, the CLLP recommends actions that lead toward achieving the subgoal. A path from the current node to the subgoal is constructed through the iterative execution of these actions. Subgoals reached within a predefined number of steps, $k$, are incorporated into the graph, while those that are not are discarded.
$\mathcal{P}:\underbrace{\mathcal{S}}_{\text{current state}}\times\underbrace{\mathcal{S}}_{\text{subgoal state}}\rightarrow\underbrace{\mathcal{A}}_{\text{action}}$
Value The value function estimates the distance from a current state to the final solution. This estimation is used to guide the selection and expansion of nodes, influencing the overall search strategy.
$\mathcal{V}:\underbrace{\mathcal{S}}_{\text{state to evaluate}}\rightarrow\underbrace{\mathbb{R}}_{\text{value of the state}}$
Behavioral Cloning Policy The policy $\Pi_{\text{BC}}$ is a decision-making function that maps the current state to an action. It encapsulates the strategy derived from the learning process, guiding the agent’s actions towards achieving the final goal.
$\Pi_{\text{BC}}:\underbrace{\mathcal{S}}_{\text{current state}}\rightarrow\underbrace{\mathcal{A}}_{\text{action}}$
D.2 Supervised Objectives
Each expert trajectory is defined as a sequence of states and corresponding actions $(s_{0},a_{0}),\ldots,(s_{n-1},a_{n-1}),s_{n}$ that delineate a path to a solution. The training methodology leverages this data through several key self-supervised imitation mappings:
  • A $k$-subgoal generator that maps a state $s_{i}$ to a future state $s_{i+k}$, simulating the achievement of intermediate goals.
  • A value function that estimates the remaining steps to the solution by mapping state $s_{i}$ to a numerical value $(i-n)$, representing the estimated distance from the goal.
  • A policy that maps each state-action pair $(s_{i},s_{i+d})$, with $d\leq k$, to the corresponding action $a_{i}$, thereby guiding the decision-making process towards the solution.
Appendix E Offline pretraining: trajectories
E.1 Rubik’s Cube
E.1.1 Random
To construct a random successful trajectory, we performed 20 random permutations on an initially solved Rubik’s Cube and took the reverse of this sequence, replacing each move with its reverse. Such solutions are usually sub-optimal since random moves are not guaranteed to increase the distance from the solution. They can even make loops in the trajectories. However, a cube scrambled with 20 moves is usually close to a random state, so such trajectories give a decent space coverage.
E.1.2 Beginner, CFOP
Beginner and CFOP are algorithms commonly used by humans. They solve the cube by ordering the stickers layer by layer. Because of that, the solutions are highly structured and long – usually between 100 and 200 moves. Both algorithms are composed of several subroutines that help building the consecutive layers. Thus, the structure of such solutions highly resembles the subgoal search.
E.1.3 Kociemba
The Kociemba two-stage solver leverages the algebraic structure of the Rubik’s Cube. In the first stage, its goal is to enter a specific subgroup. Since that subgroup is much smaller than the whole space, completing the solution may be done efficiently. Kociemba finds reasonably short solutions (usually between 20 and 40 moves) and works reasonably fast.
E.1.4 Size of datasets
For training the components on a dataset collected by a single solver, we generate 100 000100000100\,000100 000 trajectories. For the experiment with diverse experts, each solver generates 25 0002500025\,00025 000 trajectories for a total of 100 000100000100\,000100 000.
E.2 INT
Trajectories are constructed from sequences of axiom applications, similarly to [63], who followed [60]. A set of up to 15 (out of 18) axioms is first selected, and then a random axiom order is set and validated. Finally, a proof is converted to a relevant trajectory. Approximately 500,000 trajectories were generated for model pre-training.
We capped the number of axioms at 15 because some pairs of axioms (eg. terminal axioms) cannot be in one trajectory.
E.3 N-Puzzle
To collect data for N-puzzles, we utilized an algorithm that initially arranges block number $1$, followed by block number $2$, and so forth, as depicted in Figure 12. The training set comprises approximately $10,000$ trajectories.
E.4 Sokoban
To collect trajectories for Sokoban, we used a trained MCTS agent that gathered approximately $100,000$ trajectories.
Appendix F Algorithms
F.1 Best-First Search
Overview
Best-First Search greedily prioritizes node expansions with the highest heuristic estimates, aiming for paths that likely lead to the goal. While not ensuring optimality, BestFS provides a simple yet efficient strategy for navigating complex search spaces. The high-level pseudocode for BestFS is outlined in Algorithm 1, and the detailed pseudocode is presented in Algorithm 2.
Algorithm 1 Pseudocode for Best-First Search
while has nodes to expand do
   Take node $N$ with the highest value
   Select children $n_{i}$ of $N$
   Compute values $v_{i}$ for the children
   Add $(n_{i},v_{i})$ to the search tree
end while
Heuristic
In our implementation, we adhere to the Best-First Search principle by utilizing the learned value function, a common practice in the planning domain [6, 14, 63, 35]. It should be noted that in each of our experiments, all the compared algorithms use the same value function network. This way we ensure that the differences come solely from the algorithmic part.
Selecting children
When expanding a node during search, the standard BestFS algorithm adds all its children. However, in our implementation, we aimed to reduce the search tree size by selecting only the most promising children. We achieve this by sorting the children according to their probability distribution predicted by the policy network. For choosing the final subset of children, we employ two approaches. In the simpler variant, we always select the top $k$ actions. In the second variant, we add top children until their cumulative probability exceeds a fixed threshold $t_{conf}$.
This pruning does not adversely affect the standard algorithm, as nodes are still chosen based on their heuristic values, while the threshold sets a practical limit on the search space. Our results demonstrate that BestFS tends to perform much better with a confidence threshold (Figure 28). However, its performance is highly sensitive to this threshold as it balances exploration and exploitation, illustrating the impact of different confidence thresholds on success rates.
Figure 28: Comparison of success rates for the BestFS algorithm on the Rubik’s Cube with various confidence threshold values. BestFS-X represents the BestFS algorithm with the confidence threshold set to X. Left: The plot displays the achieved success rate relative to the graph size. Right: The plot illustrates the success rate for a budget of 500 nodes.
Completeness
In the Rubik’s Cube environment with random trajectories, the subgoal methods solve more instances than BestFS given a low search budget, but with more resources, BestFS takes the lead (see Figure 2). Also, in other experiments, we may observe that BestFS typically requires higher computational budget to solve the simplest instances, but its performance increases considerably with more resources.
That behavior is related to the fact that the search trees built by hierarchical methods are much sparser because the branching occurs only in the high-level nodes. On the other hand, the low-level algorithms can cover a higher fraction of the space. On the extreme, if we used all the available actions for every expansion, the low-level search would be guaranteed to find a solution if one exists. Our mechanism of selecting the actions removes that guarantee. However, at the same time, it drastically improves performance (compare BestFS-0.7 with BestFS-0.99 which is complete), which makes it a much better choice for our study.
We note that the high-level algorithms could be made complete, as proposed in [36, 63]. However, to maximize the efficiency we choose to keep the tested algorithms in their original form. The ability to search with sparse trees not only lets the methods advance fast, but also withdraw quickly if the branch does not lead to the solution (is a dead end).
Hyperparameters To identify the most suitable solving parameters, we used grid search. Initially we grid over coarse values (namely $0.1$, $0.2$, $0.3$, $0.4$,$0.5$, $0.6$, $0.7$, $0.8$, $0.9$, and $0.99$). Then we check finer values (with precision of $0.05$) around the best-performing threshold. The best-performing thresholds range from $0.6$ to $0.85$, depending on the environment and the components that are used.
For determining the best number of top actions $k$ for the simpler variant, we simply check every possible number of actions. Usually selecting $2$ actions is by far the best choice.
Details regarding hyperparameters of the networks are listed in Appendix D.1.
Algorithm 2 Complete pseudocode for Best-First Search
value function network $V$,
policy $\rho_{BFS}$
predicate of solution solved
function search($s_{0}$)
$T\leftarrow\emptyset$priority queue}
$T$.push$((V(s_{0}),s_{0}))$
[…]
seen.add$(s_{0})$seen is a set}
[…]
   _$\_,s\leftarrow T.\textsc{extractMax}()$select node with the highest value}
   $\textit{actions}\leftarrow\rho_{BFS}(s)$
   for $a\textbf{ in }\textit{actions}$ do
     $s^{\prime}\leftarrow\textsc{envStep}(s,a)$
     if s $s^{\prime}\textbf{ in }\textit{seen}$ then
        continue
     end if
     $\textit{seen}.\textsc{add}(s^{\prime})$
     $\textit{parents}[s^{\prime}]\leftarrow s$
     $T.\textsc{push}((V(s^{\prime}),s^{\prime}))$
     if $(s^{\prime})$ then
        {solution found}
        return extractLowLevelTrajectory($s^{\prime}$, parents)
     end if
   end for
end while
return $\mathtt{False}$solution not found}
F.2 Monte Carlo Tree Search
Overview
Our Monte Carlo Tree Search (MCTS) solver, designed for a single-player setting, is based on the AlphaZero framework [51]. The high-level workflow of MCTS is illustrated in Figure 29, and detailed pseudocode is provided in Algorithm 3.
The algorithm’s operation consists of four primary stages:
  • Selection: The most promising node is selected using Polynomial Upper Confidence Trees (PUCT), augmented with an exploration weight to strike a balance between exploiting known strategies and investigating new pathways.
  • Expansion: The selected node is expanded, generating new child nodes that correspond to prospective future actions. This expansion widens the search tree and enables the exploration of various outcomes.
  • Simulation: Following the AlphaZero approach [51], policy and value networks replace traditional simulations. The policy network suggests favorable moves, while the value network predicts their probability of success, directing the algorithm towards beneficial trajectories.
  • Backpropagation: The insights derived from the networks are used to update node values, improving future decision-making.
Figure 29: Schematic diagram of the MCTS algorithm in our implementation. Arrows show policy network probabilities and node values are valued network predictions. Q values, calculated via PUCT, integrate these with exploration-exploitation balance.
Hyperparameters
In the MCTS algorithm, the parameters were set as follows: sampling temperatures were chosen from [0, 0.5, 1]. The number of steps varied between 200 and 1000, and the number of simulations ranged from 5 to 300. The discount factor and exploration weight were consistently set at 1.
Algorithm 3 MCTS Solver
Number of simulations: $N_{s}$
Discount factor: $\gamma$
Exploration weight: $c_{\text{puct}}$
Sampling temperature: $\tau$
Value function: $V$
Environment model: $M$
Initial state: $initial\_state$ from env
function Search(($initial\_state$))
$root\leftarrow initial\_state$
$iteration\leftarrow 0$
while $iteration<N_{s}$ do
   $node\leftarrow root$
   while $node$ is not a leaf do
     $node\leftarrow\textsc{selectChild}(node)$, according to PUCT formula
   end while
   $leaf\leftarrow node$
   Expand the leaf using the environment model $M$, policy $\pi$, value function $V$, and discount factor $\gamma$
   Backpropagate results through the path to update $N,W,Q$
   $iteration\leftarrow iteration+1$
end while
$best\_child\leftarrow\text{Sample child of the }root\text{ according to $\tau$and $N$}$
return action leading to $best\_child$
F.3 A Search
Overview
Like Best-First Search, A* prioritizes the exploration of promising nodes. However, A* strategically guides its search by incorporating both the actual cost to reach a node and a heuristic estimate of the remaining distance to the goal. This way it balances the greedy exploitation and conservative exploration. The high-level pseudocode for A* is outlined in Algorithm 4, and the detailed pseudocode is presented in Algorithm 5.
Algorithm 4 Pseudocode for A*
while has nodes to expand do
   Take node $N$ with the highest value
   Select children $n_{i}$ of $N$
   Compute values $v_{i}$ for the children
   Compute depth $d_{i}$ for the children
   Add $(n_{i},\lambda d_{i}+v_{i})$ to the search tree
end while
Heuristic
A* guidance is achieved through the following cost function:
$f(node)=\lambda g(node)+h(node)$
where:
  • $g(node)$: The cost to reach $node$ from the start state, in our case its depth in the search tree.
  • $h(node)$: A heuristic estimate of the cost from $node$ to the goal state.
  • $\lambda$: A scaling factor balancing the influence of actual cost and heuristic estimate.
For heuristic $h$, we used a value network, like for BestFS (see Appendix F.1). If the heuristic used for A* is admissible, i.e. it never overestimates the cost of reaching the goal, A* is guaranteed to find an optimal solution. For instance, if we used $h(node)\equiv 0$, A* would reduce to the Dijkstra algorithm. The heuristic that we learn is not guaranteed to be admissible. Firstly, it estimates the distance according to the demonstrations, which is always an upper bound for the optimal distance. Secondly, the approximation errors introduce additional uncertainty. However, our main focus is on finding any solution, not necessarily an optimal one.
Selecting children
During the search, A* maintains a priority queue of nodes to be explored. Similarly to BetsFS (Appendix F.1) for reducing the search tree size, we select the most promising children. At each iteration, the node with the lowest $f(node)$ value is selected for expansion. The algorithm proceeds until the goal state is reached or the computational budget is exceeded.
Figure 30: Figures presented above illustrate the impact of depth cost scaling on the overall success rate of the A* algorithm on Sokoban, employing a confidence threshold of 0.85. In most experiments, the smaller the depth scaling factor is, the better is the final success rate. The left figure shows the success rate curves for different choices of cost weight $\lambda$, while the right plot compares those variants for a fixed budget of 500 computation nodes.
Hyperparameters
The key parameter for A* is the cost weight $\lambda$. On the extreme, setting $\lambda=0$ reduces A* to greedy BestFS, while setting $\lambda=\infty$ makes it equivalent to Breadth-First Search. By tuning that parameter, we control the trade-off between exploration and exploitation of the search.
To tune the depth parameter for our experiments, we grided over values $[0.1,0.2,0.5,1,2,5,10]$. However, usually the best choice was to keep the cost weight low ($0.1$ or $0.2$, see Figure 30). While conservative search allows A* avoid more dead-ends than BestFS (see Table 8), usually greedy steps lead to finding the solution much faster.
Algorithm 5 Complete pseudocode for A*
Initial state $initial\_state$
Heuristic function $\mathcal{V}(node)$
function AStarSearch(($initial\_state$))
$OpenSet\leftarrow\{initial\_state\}$
$CameFrom\leftarrow$ an empty map
$GCost\leftarrow$ map with default value of $\infty$
$GCost[initial\_state]\leftarrow 0$
$FCost\leftarrow$ map with default value of $\infty$
$FCost[initial\_state]\leftarrow\mathcal{V}(initial\_state)$
while $OpenSet$ is not empty do
   $current\leftarrow$ node in $OpenSet$ with the lowest $FCost$ value
   if $current=goal\_state$ then
     return $ReconstructPath(CameFrom,current)$
   end if
   $OpenSet.remove(current)$
   for each neighbor of $current$ do
     $TentativeGCost\leftarrow GCost[current]+dist\_between(current,neighbor)$
     if $TentativeGCost<GCost[neighbor]$ then
        $CameFrom[neighbor]\leftarrow current$
        $GCost[neighbor]\leftarrow TentativeGCost$
        $FCost[neighbor]\leftarrow GCost[neighbor]+\mathcal{V}(neighbor)$
        if $neighbor$ not in $OpenSet$ then
          $OpenSet.add(neighbor)$
        end if
     end if
   end for
end while
return failure
F.4 kSubS
kSubS can be seen as a special case of Adaptive Subgoal Search, but with just one generator $\rho_{k}$. For details see Algorithm 6. Both methods are introduced and studied in [14, 63].
F.5 AdaSubS
Algorithm 6 Complete pseudocode for Adaptive Subgoal Search
$C_{1}$ max number of nodes,
$V$ value function network,
$\rho_{k_{0}},\ldots,\rho_{k_{m}}$ subgoal generators,
Solved predicate of solution
function Solve(($s_{0}$))
$T\leftarrow\emptyset$priority queue with lexicographic order}
$parents\leftarrow\{\}$
for $k$ in $k_{0},\ldots,k_{m}$ do
   $T.push((k,V(s_{0})),s_{0})$
end for
$seen.add(s_{0})$$seen$ is a set}
while $0<\text{len}(T)\textbf{ and }\text{len}(seen)<C_{1}$ do
   $(k,\_),s\leftarrow T.extract\_max()$
   $subgoals\leftarrow\rho_{k}(s)$
   for $s^{\prime}$ in $subgoals$ do
     if $s^{\prime}$ not in $seen$ then
        if Is_Valid(s, s’) then
          $seen.add(s^{\prime})$
          $parents[s^{\prime}]\leftarrow s$
          for $k$ in $k_{0},\ldots,k_{m}$ do
            $T.push((k,V(s^{\prime})),s^{\prime})$
          end for
          if Solved(s’) then
            return ExtractLowLevelTrajectory(s’, parents)
          end if
        end if
     end if
   end for
end while
return $\mathtt{False}$
F.6 HIPS and HIPS-$\varepsilon$
Here we show a pseudocode for HIPS and HIPS-$\varepsilon$ methods. For details see Alg. 7
Algorithm 7 Complete pseudocode for HIPS with BestFS-PHS* and VQ-VAE
$C_{1}$ max number of nodes,
$VAE$ Variational Autoencoder for subgoal generation,
Solved predicate of solution,
$\epsilon$ exploration parameter for balancing,
$V$ value function for PHS* cost estimation
function Extended_HIPS_Solve(($s_{0}$))
Initialize search data structures, including priority queues.
$seen.add(s_{0})$Track seen states}
while search conditions are met do
   Use Best-First Search strategy to select a state $s$.
   Generate subgoals $subgoals\leftarrow VAE(s)$.
   for each $s^{\prime}$ in $subgoals$ do
     if $s^{\prime}$ not seen and is valid then
        Evaluate $s^{\prime}$ using $V$ for PHS* cost.
        Update priority queue based on PHS* cost.
        if Solved(s)$(s^{\prime})$ then
          return Construct solution path.
        end if
     end if
   end for
end while
return $\mathtt{False}$Solution not found}
Appendix G Statistical analysis of high-level and low-level algorithms
Environment Algorithm Tree size Number of leaves Branching factor Solution length Solution length (subgoals)
N-Puzzle BestFS $354.43$ $1.34$ $1.0$ $354.08$
A* $354.09$ $1.34$ $1.0$ $353.56$
MCTS $742.04$ $371.52$ $2.0$ $347.43$
kSubS-8 $353.66$ $1.0$ $1.0$ $353.66$ $45.67$
Sokoban BestFS $185.24$ $36.88$ $1.22$ $48.98$
A* $85.04$ $12.22$ $1.43$ $45.68$
MCTS $255.0$ $128.0$ $2.0$ $45.1$
kSubS-8 $101.92$ $6.6$ $1.06$ $46.88$ $7.23$
Rubik’s Cube BestFS $152.25$ $58.02$ $1.65$ $48.92$
A* $185.23$ $69.57$ $1.64$ $45.46$
MCTS $716.46$ $358.73$ $2.0$ $33.32$
kSubS-4 $303.52$ $133.44$ $1.12$ $73.58$ $26.65$
Table 3: Average values of tree size, number of leaves, branching factor (average number of children), and solution length were calculated for 100 boards solved by all presented algorithms. Additionally, for the subgoal method, the average number of subgoals on the winning path was determined.
Figure 31: The distribution of solution length in Sokoban. The right part of each plot illustrates the distribution for the methods that we used. The left part corresponds to the optimal solutions for the tested instances obtained using Breadth-First Search. These algorithms were evaluated on 494 commonly solved instances.
Figure 32: The average difference between the solutions found by each algorithm and the optimal solutions for the Sokoban environment. These algorithms were evaluated on 494 commonly solved instances.
Appendix H Hierarchical search
Figure 33: Overview of the search methods under consideration, accompanied by illustrative examples depicted in various plots for each method. Specifically, straight blue lines are utilized to represent low-level actions that occur within the search space. In contrast, long skip connections are used to symbolize subgoals within the search process.
Appendix I Further discussion on HIPS results
HIPS and HIPS-$\varepsilon$ [35, 36] are recent hierarchical search algorithms proposing to generate subgoals with variational autoencoders. We attempted to use HIPS and HIPS-$\varepsilon$ in greedy and prior-informed variations, and for all HIPS methods, the cost of inference was prohibitively high.
To compare these methods, we used A*-generated data from HIPS papers, in contrast to all other experiments (which use data generated by us).
Our evaluation, illustrated in Figure 34, shows that HIPS uses 100x more low-level nodes in search than comparable subgoal search methods and baselines - despite relatively similar subgoal efficiency as calculated in relevant papers. These findings informed our decision not to evaluate HIPS in the rest of the paper.
Figure 34: A comparison of high-level and low-level node budgets for considered methods: HIPS, subgoal search methods, and baselines on N-Puzzle. The low-level node budget represents the number of all states that have ever been visited during the search. The bimodal distribution indicates that HIPS methods use disproportionately (over 100x) more low-level nodes than comparable subgoal search methods and baselines. This directly translates to prohibitively slow solving time.